Make a sketch, roughly to scale, to find the x- and y-components of the

following vectors. In each case, find the vector’s magnitude and the

angle, measured counterclockwise, that it makes with the +x-axis.

Given. (a) Magnitude 7.40 m, at an angle 30 °

; (b) magnitude 15 km, at

an angle 225 °

; and (c) magnitude 9.30 cm, at an angle 323 °

1
Expert's answer
2021-08-29T16:56:40-0400

a)


"x=7.4\cos{30}=6.4\ m\\y=7.4\sin{30}=3.7\ m"


b)


"x=15\cos{225}=-10.6\ km\\y=15\sin{225}=-10.6\ km"


c)


"x=9.3\cos{323}=7.4\ cm\\y=9.3\sin{323}=-5.6\ cm"